OpenStudy (anonymous):

solve: (x+7)^(2/3)=9

OpenStudy (anonymous):

take (3/2) of both side

OpenStudy (anonymous):

there's availability for 2 answers though

OpenStudy (anonymous):

\[(x+7)=9^{\frac{3}{2}}\]

OpenStudy (anonymous):

yes because you don't know whether the you have 27 or -27

OpenStudy (lone001):

x+7=9^(3/2) x+7=27 x=20

OpenStudy (anonymous):

so there are to possiblities

OpenStudy (anonymous):

namely that \[x+7=27\] or \[x+7=-27\] first one gives \[x=20\] second one gives \[x=-34\]

OpenStudy (anonymous):

left hand side of your equation says the cube root of x+7 squared. if x = -34 then x + 7 = -27. if you take the cube root you get -3.. square gives you 9. that is why you have two answers

OpenStudy (anonymous):

hope that is clear